# Multiplicities of eigenvalues and quadratic representations of integers

### 16 Mar 2026 v1

@@ Abstract.
We study the set $\mathcal{M}$} of all multiplicities of non-zero eigenvalues for the Laplace operator on a two-dimensional rectangle or torus. We show that for a rectangle with the side length ratio $r$, $\mathcal{M}$$\mathbb{N}$}, the set of all positive integers, if and only if $r^{2}$ is rational. For a torus whose generating vectors have a length ratio $r$ and the angle between them $\theta$, we show that $\mathcal{M}$} is an infinite set if and only if both $r\cos\theta$ and $r^{2}$ are rational. In this case, ${\mathcal{M}}=2\mathbb{N},4\mathbb{N}$, or $6\mathbb{N}$, and we obtain a characterization for each of these cases in term of $r\cos\theta$ and $r^{2}$. In the case when at least one of $r\cos\theta$ or $r^{2}$ is irrational, we show that ${\mathcal{M}}=\{2\}$ or $\{2,4\}$, and obtain a characterization for these cases. We prove these results by studying the number of integral lattice points on dilated ellipses.
@@ 1. Introduction
In general, it is difficult to explicitly compute eigenvalues or determine their multiplicities. For planar domains, explicit formulas for the eigenvalues are known only for a few cases, such as circles, rectangles, equilateral triangles, hemi-equilateral triangles, and isosceles right triangles (see [11]). For string vibrations, all eigenvalues are simple. On a circular planar drumhead, the first eigenvalue is simple while all other eigenvalues have multiplicities $2$. For the unit ball in ${\mathbb{R}}^{n}$, $n\geq 2$, the set of multiplicities of eigenvalues for the Dirichlet Laplacian is identical to the set of all dimensions of the spaces of spherical harmonics of degree $k$ ($k\geq 0$), which is given by ${\mathcal{M}}(B_{n})=\{1,n,C(k+n-1,n-1)-C(k+n-3,n-1)\mid k\geq 2\}$, where $C(\cdot,\cdot)$ denotes the combination number. The set of multiplicities for the Laplace operator on the unit sphere in ${\mathbb{R}}^{n}$ is the same as that of the unit ball (see [16, Chapter IV] for relevant material on spherical harmonics).
In [7], the authors studied the multiplicities of eigenvalues for the Laplace operator on rectangles. They observed that on the rectangle $\Omega_{a,b}=[0,a]\times[0,b]$, ${\mathcal{M}}=\{1\}$ if $(a/b)^{2}\notin\mathbb{Q}$, and showed that squares satisfy property ($M$), as a consequence of Legendre’s formula on the number of ways an integer can be represented as a sum of squares of integers. (A domain in ${\mathbb{R}}^{n}$ is said to satisfy property ($M$) if the set of multiplicities of eigenvalues for the Dirichlet Laplacian is $\mathbb{N}$} [7].) In [9], Heimrath obtained formulas for multiplicities of eigenvalues on rectangles such that $(a/b)^{2}\in\{2,3,7\}$ through generalizing Legendre’s formula and as a consequence, showed that these rectangles satisfy property ($M$). The problem of whether or not $\Omega_{a,b}$ satisfying property ($M$) is characterized by $(a/b)^{2}\in{\mathbb{Q}}$ was left open in [7] and [9].
The first result in this paper answers this question affirmatively:
@@ 4-4 $\pm I,\pm\big(\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\big)$
This results in two distinct values for $x^{\prime 2}$: $x^{\prime 2}=x^{2}$ or $x^{\prime 2}=cy^{2}$.
If $c$ is not the square of a rational number, then for any $y\neq 0$, the value $y\sqrt{c}$ is irrational. Consequently, the second root yields no integer solutions. The only integer solutions are $(x,y)$ and (x,y)(-x,-y). In this case $R_{\tilde{f}}(\mathbb{R^{+}})=\{0,2\}$.
If $c=r^{2}$ for some $r\in\mathbb{Q}$, then $x^{\prime}=\pm yr$ is an integer provided $yr$ is an integer. If $x\neq yr$, then the first and the second roots provide distinct values for $x^{\prime 2}$. Each value of $x^{\prime 2}$ produces two values for $x^{\prime}$, and since $y^{\prime}=L/x^{\prime}$, the pair $(x^{\prime},y^{\prime})$ is uniquely determined. Thus in this case $N(z)=4$. ∎

### 02 Apr 2026 v2

@@ Abstract.
We study the set $\mathcal{M}$} of all multiplicities of non-zero eigenvalues for the Laplace operator on a two-dimensional rectangle or torus. We show that for a rectangle with the side length ratio $r$, $\mathcal{M}$$\mathbb{N}$}, the set of all positive integers, if and only if $r^{2}$ is rational. For a torus whose generating vectors have a length ratio $r$ and the angle between them $\theta$, we show that $\mathcal{M}$} is an infinite set if and only if both $r\cos\theta$ and $r^{2}$ are rational. In this case, ${\mathcal{M}}=2\mathbb{N},4\mathbb{N}$, or $6\mathbb{N}$, and we obtain a characterization for each of these cases in terms of $r\cos\theta$ and $r^{2}$. In the case when at least one of $r\cos\theta$ or $r^{2}$ is irrational, we show that ${\mathcal{M}}=\{2\}$ or $\{2,4\}$, and obtain a characterization for these cases. We prove these results by studying the number of integral lattice points on dilated ellipses.
@@ 1. Introduction
In general, it is difficult to explicitly compute eigenvalues or determine their multiplicities. For string vibrations, all eigenvalues are simple. On a circular planar drumhead, the first eigenvalue is simple while all other eigenvalues have multiplicities $2$. For the unit ball in ${\mathbb{R}}^{n}$, $n\geq 2$, the set of multiplicities of eigenvalues for the Dirichlet Laplacian is identical to the set of all dimensions of the spaces of spherical harmonics of degree $k$ ($k\geq 0$), which is given by ${\mathcal{M}}(B_{n})=\{1,n,C(k+n-1,n-1)-C(k+n-3,n-1)\mid k\geq 2\}$, where $C(\cdot,\cdot)$ denotes the combination number. The set of multiplicities for the Laplace operator on the unit sphere in ${\mathbb{R}}^{n}$ is the same as that of the unit ball (see [17, Chapter IV] for relevant material on spherical harmonics).
For planar domains, explicit formulas for the eigenvalues are known only for a few cases, such as circles, rectangles, equilateral triangles, hemi-equilateral triangles, and isosceles right triangles (see [11]). In [15], Pinsky obtained a formula for multiplicities of eigenvalues on the equilateral triangles and show that the set of multiplicities of eigenvalues is $\mathbb{N}$. In [7], the authors studied the set of multiplicities of eigenvalues for the Laplace operator on rectangles. They observed that on the rectangle $\Omega_{a,b}=[0,a]\times[0,b]$, ${\mathcal{M}}=\{1\}$ if $(a/b)^{2}\notin\mathbb{Q}$, and showed that squares also satisfy property ($M$), as a simple consequence of Legendre’s formula on the number of ways an integer can be represented as a sum of squares of integers. (A domain is said to satisfy property ($M$) if the set of multiplicities of eigenvalues for the Dirichlet Laplacian is $\mathbb{N}$} [7].) In [9], Heimrath obtained formulas for multiplicities of eigenvalues on rectangles such that $(a/b)^{2}\in\{2,3,7\}$ through generalizing Legendre’s formula and as a consequence, showed that these rectangles satisfy property ($M$). The problem of whether or not $\Omega_{a,b}$ satisfying property ($M$) is characterized by $(a/b)^{2}\in{\mathbb{Q}}$ was left open in [7] and [9]. The first result in this paper answers this question affirmatively:
@@ 4-4 $\pm I,\pm\big(\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\big)$
Thus $x^{\prime 2}=x^{2}$ or $x^{\prime 2}=cy^{2}$.
If $c$ is not the square of a rational number, then the only integer solutions are $(x,y)$ and (x,y)(-x,-y). Therefore, in this case, $R_{\tilde{f}}(\mathbb{R^{+}})=\{0,2\}$.
If $c=r^{2}$ for some $r\in\mathbb{Q}$, then the equation $z=f(x,y)$ has 4 distinct integer solutions:
$\pm(x,\;y)\quad\text{and}\quad\pm(ry,\;x/r),$
provided both $ry$ and $x/r$ are integers, and $x\not=ry$. Moreover, the number of solutions is 2 if $x=ry$. Thus, in this case, $R_{\tilde{f}}(\mathbb{R}^{+})=\{0,2,4\}$. ∎
@@
Remark.
It would be of interest to generalize Theorem 1.2 and Theorem 1.3 to higher dimensions. Both theorems suggest that a domain whose set of multiplicities is infinite is the exception rather than the norm. The collection of rectangles or tori with such a property has zero measure among their peers. It would be interesting to quantify and understand this phenomenon in general.